# When can isotropy help adapt LLMs’ next word prediction to numerical domains?

### v1

@@ 1 Introduction
Our key contributions include: (i) We consider a log-linear model for LLMs and demonstrate theoretically why hidden representations must exhibit structure to address the shift-invariance problem of the softmax function. (ii) We take a deeper look into the hidden representations of pre-trained models and show how isotropy preserves the structural integrity of representations. In particular, we derive an upper bound for the Jacobian matrix which collects all first-order partial derivatives of self-attention with respect to the input pattern and show that $m$ largest eigenvectors of the LLM hidden representations minimize the gradient norm of self-attention. Then by projecting the representations into lower dimensions using these $m$ largest eigenvectors, we find the isotropy within the clusters in the contextual embedding space. (iii) Finally, we provide a comprehensive evaluation across $12$ real and $10$ synthetic
@@
time series datasets over $6$ different LLMs.
@@ 4 Study of isotropy in LLM hidden representations
Analysis settings. For illustrative purposes, in this section, we present two examples from a specific numerical domain, such as transport, and for a specific language model (i.e., GPT-2) that illustrate the conditions under which isotropy is preserved in LLM representations.
@@
In Section 5, we provide comprehensive evaluations across $22$ different datasets from various numerical domains and other $5$ different language models with different architectures and tokenization techniques. For this specific analysis, we select two datasets from transport as they are dynamic, noisy, time-varying, and thus hold all primary characteristics of the numerical data across various domains. Specifically, we use two ways of sharing signals which are typically used for signal transmission in transportation settings, as shown in Figure 2. We call these two datasets as “Dataset $1$” and “Dataset $2$”. The downstream task here is to predict the transport signalling property over wireless channel using LLM, where Dataset $1$ causes good downstream performance (i.e., near optimal signal prediction), while the Dataset $2$ causes bad downstream task performance (i.e., high error in signal prediction). We use NMSE as a performance metric for the numeric downstream task because it is widely used for signal prediction. We deploy the first six layers of GPT2 Radford et al. (2019) and use the datasets and simulation setups from Liu et al. (2024), which are the standard settings for time series forecasting. We predict $L=4$ future signaling properties based on the historical $T=16$ signal properties through time series forecasting using GPT2. The training and validation dataset contains $8,000$ and $1,000$ samples, respectively, with user velocities uniformly distributed between $10$ km/hour and $100$ km/hour. The test dataset contains ten velocities ranging from $10$ km/hour to $100$ km/hour, with $1,024$ samples for each velocity.
@@ 5 Experiments
Datasets. We conduct a comprehensive evaluation using $12$ different real time series datasets from various numerical domains, including energy, nature, finance, healthcare, retail and transportation (data sources can be found in Table 4 of Appendix E). We also illustrate our findings using KernelSynth Ansari et al. (2024) (see Algorithm 1 in Appendix E for details), a method that generates $10$ more synthetic datasets via Gaussian processes in Section 5. In Section 4.1, we show how the LLM hidden representation varies for two different datasets from the same numerical domain and its impact on downstream task performance. So we select two different datasets from each numerical domain (for good and bad downstream task performance examples) for qualitative and quantitative analysis as shown in Table 3.
@@ 5.1 Qualitative Analysis
In this section, we analyze the time series forecasting by the baseline LLMs qualitatively. We focus on synthetically generated time series for a controlled analysis of different types of time series patterns which belong to $5$ different domains, such as linear, seasonality, trend, non-linear and Gaussian perturbation.
We are particularly interested in the isotropic measurement in the LLM’s last layer as it is related to the logits and probabilistic inference. So all isotropic measure provided in this section is based on the last layer of the baselines.
We begin by analyzing time series forecasting performance (i.e., NMSE) for different baselines and its relation with isotropy in Figure 6. For instance, in Figure 6 b, the inter-type cosine similarity in hidden embedding spaces of Chronos-T5 for seasonality (Dataset 1) is $-0.00007$ and for seasonality (Dataset 2) is $0.0047$. This indicates that stronger isotropy exists (i.e., inter-type cosine similarity value is close to 00) for Chronos-T5 for seasonality (Dataset 1) which preserves the structure in the Chronos-T5’s hidden representations and causes good downstream task performance. On the other hand, a weaker isotropy exists (i.e., inter-type cosine similarity value is far from 00) for Chronos-T5 in seasonality (Dataset 2), causing a lack of structure in Chronos-T5’s hidden representations, and thus leading to bad downstream task performance. Moreover, the NMSE for seasonality (Dataset 1) is lower for Chronos-T5, PatchTST and Lag-Llma, and higher for Chronos-Bolt and Morai, while the NMSE is higher for all the baselines except Morai for non-linear (Dataset 2). A similar analysis can also be observed for all other synthetic datasets and baselines in Figure 6. This shows that the same dataset from any particular domain may cause different forecasting performances for different baselines, as it generates different contextual embedding spaces for language models with different architectures and tokenization techniques (See Appendix F for full illustration).
Next, we examine the influence of isotropy on forecasting performance in two important scenarios: a) different input context lengths, and b) different levels of noises in the input data. The first scenario is important as it provides an analysis on selecting proper input context length rather than selecting the length through random trials and errors. The second scenario is important as it gives us ideas on how the level of noise in noisy data impacts performance, since the data in the real world is mostly noisy.
Isotropy in different input context lengths. We first analyze the impacts of isotropy under varying input context lengths. We select Choronos-T5 as an example model and non-linear (Dataset 1) as an example dataset. In Figure 11 in Appendix F, we show how the hidden representations of Chronos-T5 vary for two different input context lengths, such as $L=500$ and $L=500$, for non-linear (Dataset 1), which gives an indication of having
@@
different isotropic measures for different context lengths.
In Figure 7, we compare the NMSE vs isotropy across two different input context lengths, $L=500$ and $L=100$, for different synthetic datasets. As can be seen from the figure, the isotropy values vary across different input context lengths and datasets. For instance, in seasonality (Dataset 1), we have (NMSE$=0.0000066$, cosine similarity$=-0.00076$) and (NMSE$=0.0793$, cosine similarity$=0.0011$) for $L=500$ and $L=100$, respectively. The decrease in isotropy significantly increases the NMSE for the input context length $L=100$. In contrast, in Linear (Dataset 2), the isotropy increases for the input context length $L=100$, which causes the decreases in NMSE for chornos-T5. In practice, the input context length is often selected randomly or through trial and error, which may cause higher forecasting errors for different datasets. Isotropy analysis enables us to understand how varying input context lengths influence the hidden representations of the language model. This insight helps guide improvements in forecasting performance by examining the isotropic properties of the contextual embedding space.
Isotropy in varying noise levels in datasets. Next, we focus on the second scenario to see the impact of noisy datasets on LLM’s performance. Again, we use the Chornos-T5 as an example language model. In Figure 8, we compare the NMSE vs isotropy across two different noise levels, one without noise, and the other with Gaussian noise with $\sigma=0.05$ standard deviation. From Figure 8, we can see consistently lower isotropy (i.e., inter-type cosine similarity far from 00) for all noisy synthetic datasets as compared to the datasets without noise. For instance, in trend (Dataset 2),
@@
we have (NMSE$=0.000024$, cosine similarity$=-0.00022$) and (NMSE$=0.0012$, cosine similarity$=0.0040$) for $\sigma=0$ and $\sigma=0.05$, respectively. The decrease in isotropy significantly increases the NMSE for the noisy dataset. In practice, the environments of many real numerical domains, such as nature and energy, are noisy and dynamic. In such an environment, it is not always possible to measure the noise in real time and take the necessary steps to clean the input time series for better performance. However, it is always possible to measure the isotropy from LLM hidden representations, which can be used as a measure of noise in the input datasets, and thus, help to improve the forecasting performance.
@@ 5.2 Quantitative Analysis
In Figure 9, we analyze the time series forecasting performance for 12 real datasets with different baselines and its relation with isotropy. For instance, in Figure 9 e, the inter-type cosine similarity in hidden embedding spaces of Morai for retail (Dataset 1) is $0.002$ and for retail (Dataset 2) is $0.1931$. This indicates that stronger isotropy exists for Morai for retail (Dataset 1) which preserves the structure in the Morai’s hidden representations and causes good downstream task performance. On the other hand, a weaker isotropy exists for Morai in retail (Dataset 2),
@@
causing a lack of structure in Morai’s hidden representations, and thus leading to bad downstream task performance. Moreover, the NMSE for retail (Dataset 2) is lower for all baselines except the Morai as compared to retail (Dataset 1), while the NMSE for energy (Dataset 1) is lower for all baselines except Lag-Llma as compared to energy (Dataset 2). A similar analysis can also be observed for all other real datasets and baselines in Figure 9. This again shows that a dataset may have different impacts on forecasting performances for different baselines, and isotropy can be used as a measure of how the dataset generates different contextual embedding spaces for different language models based on their architectures and tokenization techniques.
Finally, in Figure 10, we compare the NMSE vs isotropy for varying input context lengths to observe its impact on the real datasets. We select Lag-Llma as our example model. We compare the results for two different input context length: 1) the recommended input context length $L=144$ and the reduced input context length $L=96$. As can be seen from the figure, the violation in recommended input context length
@@
by reducing it from $L=144$ to $L=96$ not only decreases the NMSE performances, but also increases for some datasets. For instance, the inter-type cosine similarity values become close to 00, i.e., from $0.1091$ to $0.0012$ and from $0.2014$ to $0.0396$, respectively, for nature (Dataset 2) and finance (Dataset 1), which in turn improves the NMSE performances.
@@ 6 Conclusion and Limitations
$=|\mathbb{\Lambda}|_{2}\sum_{i=1}^{|\mathcal{V}|}\left(p_{i,i}+\frac{1}{2}\right)|x_{i}-X^{\top}p_{i,:}|^{2}+|\mathcal{V}|+\Delta,$
where $\Delta=|\mathbb{\Lambda}|_{2}\sum_{i\neq j}^{|\mathcal{V}|}p_{i,j}\left|\mathbb{\psi}_{j}-\sum_{q=1}^{|\mathcal{V}|}p_{i,q}\mathbb{\psi}_{q}\right|^{2}+\frac{|\mathbb{\Lambda}|_{2}}{2}\sum_{j=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}|^{2}.$
@@ Appendix F Full Visualization of PCA plots for different models
Figure 11: Variations in Chronos-T5’s hidden representations for different input context lengths for the same synthetic dataset “non-linear (Dataset 1)” : (a) Contextual embedding space for input context length $L=500$. (b) Contextual embedding space for input context length $L=100$.
F.1 Synthetic datasets

### v3

@@ 1 Introduction
Our key contributions include: (i) We consider a log-linear model for LLMs and demonstrate theoretically why hidden representations must exhibit structure to address the shift-invariance problem of the softmax function. (ii) We take a deeper look into the hidden representations of pre-trained models and show how isotropy preserves the structural integrity of representations. In particular, we derive an upper bound for the Jacobian matrix which collects all first-order partial derivatives of self-attention with respect to the input pattern and show that $m$ largest eigenvectors of the LLM hidden representations minimize the gradient norm of self-attention. Then by projecting the representations into lower dimensions using these $m$ largest eigenvectors, we find the isotropy within the clusters in the contextual embedding space. (iii) Finally, we provide a comprehensive evaluation across $12$ real and $10$ synthetic time series datasets over $6$ different LLMs. Our experiments
@@
demonstrate that different characteristics of numerical data and model architectures could have different impacts on the isotropy.
@@ 4 Study of isotropy in LLM hidden representations
Analysis settings. For illustrative purposes, in this section, we present two examples from a specific numerical domain, such as transport, and for a specific language model (i.e., GPT-2) that illustrate the conditions under which isotropy is preserved
@@
in LLM representations. In Section 5, we provide comprehensive evaluations across $22$ different datasets from various numerical domains and other $5$ different language models with different architectures and tokenization techniques. For this specific analysis, we select two datasets from transport as they are dynamic, noisy, time-varying, and thus hold all primary characteristics of the numerical data across various domains. Specifically, we use two ways of sharing signals which are typically used for signal transmission in transportation settings, as shown in Figure 2. We call these two datasets as “Dataset $1$” and “Dataset $2$”. The downstream task here is to predict the transport signalling property over wireless channel using LLM, where Dataset $1$ causes good downstream performance (i.e., near optimal signal prediction), while the Dataset $2$ causes bad downstream task performance (i.e., high error in signal prediction). We use NMSE as a performance metric for the numeric downstream task because it is widely used for signal prediction. We deploy the first six layers of GPT2 Radford et al. (2019) and use the datasets and simulation setups from Liu et al. (2024), which are the standard settings for time series forecasting. We predict $L=4$ future signaling properties based on the historical $T=16$ signal properties through time series forecasting using GPT2. The training and validation dataset contains $8,000$ and $1,000$ samples, respectively, with user velocities uniformly distributed between $10$ km/hour and $100$ km/hour. The test dataset contains ten velocities ranging from $10$ km/hour to $100$ km/hour, with $1,024$ samples for each velocity.
@@ 5 Experiments
Datasets. We conduct a comprehensive evaluation using $12$ different real time series datasets from various numerical domains, including energy, nature, finance, healthcare, retail and transportation. The sources of these open-source datasets along with their descriptions, including how each dataset is used across different LLM can be found in Table 4 of Appendix E. We also illustrate our findings using KernelSynth Ansari et al. (2024) (see Algorithm 1 in Appendix E for details), a method that generates $10$ more synthetic datasets via Gaussian processes in Section 5. As in Section 4.1, we select two different datasets from each numerical domain (as shown in Table 3) and then perform qualitative analysis with synthetic datasets and quantitative analysis with real datasets. The results of both of this analysis are provided in the next two sections.
@@ 5.1 Qualitative Analysis
In this section, we analyze the time series forecasting by the baseline LLMs qualitatively. We focus on synthetically generated time series for a controlled analysis of different types of time series patterns which belong to $5$ different domains, such as linear, seasonality, trend, non-linear and stochastic. We are particularly interested in the isotropic measurement in the LLM’s last layer as it is related to the logits and probabilistic inference as explained in Section 2. So all isotropic measure provided in this section is based on the last layer of the baselines.
We begin by analyzing time series forecasting performance (i.e., NMSE) for different baselines and its relation with isotropy in Figure 6. For instance, in Figure 6 b, we have (NMSE = $0.0000066$ and cosine similarity = $|-0.00076|$) for seasonality-Dataset 1 and (NMSE = $0.00012$ and cosine similarity = $0.0047$) for seasonality-Dataset 2 for Chronos-T5. This shows stronger isotropy exists (i.e., inter-type cosine similarity value is close to 00) in Chronos-T5’s embedding space for seasonality-Dataset 1 which preserves the structure in its hidden representations and causes good downstream task performance. On the other hand, a weaker isotropy exists (i.e., inter-type cosine similarity value is far from 00) in Chronos-T5’s embedding space for seasonality-Dataset 2, and thus causes a lack of structure in its hidden representations which leads to bad downstream task performance as compared to seasonality-Dataset 1. The NMSE and inter type cosine similarity may also vary for different language models and datasets. For example, in Figure 6 c, we can see that the the NMSE for trend-Dataset 1 is lower for PatchTST and Morai, while higher for Chronos-T5, Chronos-Bolt and Lag-Llma. In contrast, the NMSE for trend-Dataset 2 is lower for Chronos-T5, Chronos-Bolt, and Lag-Llma, while higher for PatchTST and Morai. A similar analysis can also be observed for other synthetic datasets and baselines in Figure 6 b, Figure 6 d, and Figure 6 e. This shows that any dataset from any particular domain may cause different forecasting performances for different baselines,
@@
as it generates different hidden representations in contextual embedding spaces (see Appendix F for full visualization) for language models with different architectures and tokenization techniques.
Next, we examine the influence of isotropy on forecasting performance in two important scenarios: a) different input context lengths, and b) different levels of noises in the input data. The first scenario is important as it provides an analysis that helps guide in selecting reasonable input context lengths rather than selecting the length through random trials and errors. The second scenario is important as it gives us ideas on how the level of noise in noisy data impacts performance, since the data in the real world is mostly noisy.
Isotropy in different input context lengths. We first analyze the impacts of isotropy under varying input context lengths. In Figure 7, we compare the NMSE vs isotropy across two different input context lengths, $L=500$ and $L=100$, for different synthetic datasets. We use Choronos-T5 as an example model for this experiment. As can be seen from the figure, the isotropy values vary across different input context lengths and datasets. For instance, in seasonality-Dataset 1, we have (NMSE$=0.0000066$, cosine similarity$=|-0.00076|$) and (NMSE$=0.0793$, cosine similarity$=0.0011$) for $L=500$ and $L=100$, respectively. The decrease in isotropy significantly increases the NMSE for the input context length $L=100$. In contrast, in linear-Dataset 2, we have (NMSE$=0.000025$, cosine similarity$=0.2474$) and (NMSE$=0.000009$, cosine similarity$=0.0644$) for $L=500$ and $L=100$, respectively. In this scenario, the isotropy increases for the input context length $L=100$, which causes the decrease in NMSE for chornos-T5. In Figure 11 in Appendix F, we show how the hidden representations of Chronos-T5 vary for two different input context lengths, such as $L=500$ and $L=500$, for non-linear-Dataset 1, which generates different isotropic measures for different input context lengths. In practice, the input context length is often selected randomly or through trial and error, which may cause higher forecasting errors for different datasets. Isotropy analysis enables us to understand how varying input context lengths influence the hidden representations of the language model. This insight helps guide improvements in forecasting performance by examining the isotropic properties of the contextual embedding space.
@@
Isotropy in varying noise levels in datasets. Next, we focus on the second scenario to see the impact of noisy datasets on LLM’s performance. Again, we use the Chornos-T5 as an example language model. In Figure 8, we compare the NMSE vs isotropy across two different noise levels, one without noise, and the other with Gaussian noise with $\sigma=0.05$ standard deviation. From Figure 8, we can see consistently lower isotropy (i.e., inter-type cosine similarity far from 00) for all noisy synthetic datasets as compared to the datasets without noise. For instance, in trend-Dataset 2, we have (NMSE$=0.000024$, cosine similarity$=|-0.00022|$) and (NMSE$=0.0012$, cosine similarity$=0.0040$) for $\sigma=0$ and $\sigma=0.05$, respectively. The decrease in isotropy significantly increases the NMSE for the noisy dataset. In practice, many real-world numerical domains—such as those in nature and energy—exhibit noisy and dynamic behavior. In these environments, it is often infeasible to measure noise in real time or to pre-process the input time series for improved performance. However, the isotropy in LLMs hidden representations can be readily measured, and thus, can be leveraged to enhance forecasting performance by identifying and mitigating the effects of noisy inputs in contextual embedding space.
@@ 5.2 Quantitative Analysis
In this section, we present our main results on $12$ real datasets which belong to $6$ different numerical domains such as energy, nature, finance, healthcare, retail and transportation. As our qualitative analysis in Section 5.1, we select two different datasets from each numerical domain and the isotropy measure from LLM’s last layer to show the impact of isotropy on NMSE performance for different language models.
@@
In Figure 9, we analyze the time series forecasting performance of different baselines and its relation with isotropy for different real datasets. For instance, in Figure 9 e, we have (NMSE = $0.0061$ and cosine similarity = $0.0020$) for retail-Dataset 1 and (NMSE = $0.1255$ and cosine similarity = $0.1931$) for retail-Dataset 2 for Morai. This indicates the existence of stronger isotropy in Morai’s embedding space for retail-Dataset 1 which preserves the structure in its hidden representations and causes good downstream task performance. On the other hand, a weaker isotropy exists in Morai’s embedding space for retail-Dataset 2, and thus causes a lack of structure in its hidden representations which leads to bad downstream task performance as compared to retail-Dataset 1. The NMSE and inter type cosine similarity may also vary for different real datasets and language models. For example, in Figure 9 a, we can see that the NMSE for energy-Dataset 1 is lower for Chronos-T5, Chronos-Bolt, PatchTST and Morai, while higher Lag-Llma. In contrast, the NMSE for energy-Dataset 2 is lower for Morai while higher for other baselines. A similar analysis can also be observed for other synthetic datasets and baselines in Figure 9 b and Figure 9 c. This again shows that datasets from the same numerical domain can cause varying forecasting performance across different baselines, as they generate distinct hidden representations in contextual embedding spaces depending on the language model architecture and tokenization strategy.
@@
Finally, in Figure 10, we compare the NMSE vs isotropy for varying input context lengths to observe its impact on the real datasets. We select Lag-Llma as our example model. We compare the results for two different input context lengths: 1) the recommended input context length $L=144$ and the reduced input context length $L=96$. As can be seen from the figure, the inter-type cosine similarity values become far from 00, i.e., from $0.0097$ to $0.0220$ for energy-Dataset 1 and from $0.0026$ to $0.0103$ for transport-Dataset 1, which in turn decreases the NMSE performances. On the other hand, the inter-type cosine similarity values become close to 00, i.e., from $0.1091$ to $0.0112$ for nature-Dataset 2 and from $0.2014$ to $0.0133$ for finance-Dataset 1, which in turn improves the NMSE performances. Thus, the variation in the recommended input context length may not only decrease the NMSE performances, but can also increase for some datasets.
@@ Appendix F Full Visualization of PCA plots for different models
Figure 11: Variations in Chronos-T5’s hidden representations for different input context lengths for the same synthetic dataset non-linear-Dataset 1 : (a) Contextual embedding space for input context length $L=500$. (b) Contextual embedding space for input context length $L=100$.
The full visualization of PCA plots of different models is provided below. We use the synthetic Dataset 1, and Dataset 2 from non-linear domain for illustration.